\begin{thebibliography}{99}

\bibitem{Bertram}
E.A. Bertram, 
On large cyclic subgroups of finite groups, 
{\em Proc. Amer. Math. Soc.} 56 (1976), 63--66.
\\[-16pt]

\bibitem{Magma}
W.\ Bosma, J.\ Cannon and C.\ Playoust,
The {\sc Magma} Algebra System I: The User Language,
{\em J. Symbolic Computation\/} 24 (1997), 235--265. 
\\[-16pt]

\bibitem{CoxeterMoser}
H.S.M. Coxeter and W.O.J. Moser, 
{\em Generators and Relations for Discrete Groups}, 4th Ed., 
Springer-Verlag, Berlin, 1984.
\\[-16pt]

\bibitem{Kubilius} 
J. Kubilius, 
{\em Probabilistic Methods in the Theory of Numbers}, 
Transl. Math. Monogr., vol. 11, 
American Math, Society, Providence, RI, 1964. xviii+182 pp.
\\[-16pt]

\bibitem{Kubilyus} 
I.P. Kubilyus, 
Probabilistic methods in the theory of numbers (Russian). 
{\em Uspekhi Mat. Nauk.} 11 (1956), 31--66. 
\\[-16pt]

\bibitem{Larsen} 
M. Larsen, 
How often is $84(g-1)$ achieved?, 
{\em Israel J. Math.} 126 (2001), 1--16. 
\\[-16pt]

\bibitem{Malcev} 
A.I. Mal'cev, 
On isomorphic matrix representations of infinite groups of matrices (Russian), 
{\em  Mat. Sb.} 8 (1940), 405--422; 
{\em Amer. Math. Soc. Transl.} (2) 45 (1965), 1--18. 
\\[-16pt]

\bibitem{MZ}
C.L. May and J. Zimmerman, 
The symmetric genus of large odd-order groups, 
{\em Houston J. Math\/}.44 (2018), 1--19.
\\[-16pt]

\bibitem{Narkiewicz}
W. Narkiewicz, 
{\em Number Theory}, 
World Scientific Publishing Co., Singapore, 1983. xii+371 pp.
\\[-16pt]

\bibitem{Neumann}
P.M. Neumann, 
The SQ-universality of some finitely presented groups, 
{\em J. Austral. Math. Soc.} 16 (1973), 1--6.
\\[-16pt]

\bibitem{Tao}
T. Tao, {\em What's new} 254A, Notes 9 -- second moment and entropy methods, November 2019, \\ 
\url{https://terrytao.wordpress.com/2019/11/12/254a-notes-9-second-moment-and-entropy-methods}. 
\\[-16pt]
 
\bibitem{Tenenbaum}
G. Tenenbaum, 
{\em Introduction to Analytic and Probabilistic Number Theory}, 
Grad. Stud. Math., vol. 163, 
American Mathematical Society, Providence, RI, 2015, xxiv+629 pp.
\\[-16pt]

\bibitem{Tucker}
T.W. Tucker, 
The density of the orders of triangle group quotients, 
{\em unpublished notes}, 2022. 
\\[-16pt]

\bibitem{Turan}
P. Tur{\' a}n, 
On a theorem of Hardy and Ramanujan, 
{\em J. London Math. Soc.} 9 (1934), 274--276.
\\[-16pt]

\bibitem{MQ} 
D. Young, 
Density of numbers with a large prime factor in specified arithmetic progression, 
Question on {\tt math}{\bf\em overflow}, June 2024.

\end{thebibliography}

\documentclass{article}
\usepackage{mathtools,amssymb}
\usepackage{amsthm}
\usepackage{amsmath}
\usepackage{hyperref}
\usepackage{comment}
\usepackage[left=3cm, right=3cm]{geometry}
\usepackage[shortlabels]{enumitem}
\newtheorem{theorem}{Theorem}

\newtheorem{definition}{Definition}
\newtheorem{lemma}{Lemma}
\newtheorem{proposition}{Proposition}

\newtheorem{corollary}{Corollary}
\newtheorem*{theorem*}{Theorem}
\newtheorem*{observation*}{Observation}
\newtheorem*{question*}{Question}
\newtheorem*{lemma*}{Lemma}

\DeclareMathOperator{\N}{\mathbb N}
\DeclareMathOperator{\Z}{\mathbb Z}
\DeclareMathOperator{\R}{\mathbb R}

\DeclareMathOperator{\C}{\mathbb C}

\DeclareMathOperator\Aut{Aut}

\newcommand{\Px}{\mathcal{P}_x}
\title{\Large \bf Quotient order density of ordinary triangle groups }
\author{Darius Young}
\date{9 August 2024}

\begin{document}

\maketitle
\begin{abstract}
For positive integers $r$, $s$ and $t$, the ordinary triangle group $\Delta^+(r,s,t)$ is the 
group with presentation $\langle\, x,y,z \;|\; x^r = y^s = z^t = xyz =1 \,\rangle$, 
and is infinite if and only if $1/r+1/s+1/t \le 1$. 
In this paper, it is shown that the natural density (among the positive integers) 
of the orders of the finite quotients of every such group is zero, 
taking inspiration from a 1976 theorem of Bertram on large cyclic subgroups of finite groups, 
and the Turan-Kubilius inequality from asymptotic number theory. 
This answers a challenging question raised by Tucker, based on some work 
for special cases by May and Zimmerman, and himself. 
\end{abstract}

\section{Introduction}
\label{sec:Intro}

The purpose of this paper is to prove that the natural density
of the orders of the finite quotients of every infinite ordinary hyperbolic triangle group is zero.

For positive integers $r$, $s$ and $t$, the ordinary triangle group $\Delta^+(r,s,t)$ is the 
group with presentation
$$
\Delta^+(r,s,t) = \langle\, x,y,z \;|\; x^r = y^s = z^t = xyz = 1 \,\rangle, 
$$
or more simply, $\langle\, x,y \;|\; x^r = y^s = (xy)^t = 1 \,\rangle$. 
This is a Fuchsian group and hence plays a role in the study of group actions on algebraic curves, compact Riemann surfaces 
and regular maps on surfaces, and is said to be  {\em spherical}, {\em Euclidean\/} (or {\em toroidal\/}), 
or {\em hyperbolic}, depending on whether $\frac{1}{r}+\frac{1}{s}+\frac{1}{t}$ is greater than $1$, 
equal to $1$ or less than $1$, respectively. 

In the spherical case, every such group is finite (and is cyclic or dihedral, or isomorphic to one 
of $A_4$, $S_4$ and $A_5$), while in the Euclidean case, the triangle groups are infinite but soluble, 
and in the hyperbolic case they are infinite but insoluble.  
Moreover, every hyperbolic triangle group $\Delta^+ = \Delta^+(r,s,t)$ is residually finite 
(by a theorem of Mal'cev \cite{Malcev} on Fuchsian groups), and also `SQ-universal', 
meaning that every countable group is isomorphic to a subgroup of some quotient of $\Delta^+$; see \cite{Neumann}. 

Finite quotients of triangle groups have been of great interest in many parts of mathematics for quite some time, 
and not only in group theory but also in complex analysis, geometry, graph theory and topology, dating back 
to the 1893 theorem of Hurwitz on the largest number of conformal automorphisms of a compact 
Riemann surface of genus $g > 1$, namely $84(g-1)$, and other work around that time by various mathematicians including Burnside, Dyck and Klein. 

Of particular interest are the so-called {\em smooth\/} quotients, which are those in which the 
orders of the group generators $x$, $y$ and $z$ ($= (xy)^{-1}$) are preserved, or equivalently, 
those obtained as images of a homomorphism with torsion-free kernel.
While we will not often mention smooth quotients explicitly, we note that the main theorem of this paper also holds for smooth finite quotients, since those form a subset of all quotients.

Closely related to each ordinary triangle group $\Delta^+(r,s,t)$ is the full triangle group 
$$
\Delta(r,s,t) = \langle\, a,b,c \;|\; a^2 = b^2 = c^2 = (ab)^r = (bc)^s = (ac)^t = 1 \,\rangle, 
$$
which clearly contains $\Delta^+(r,s,t)$ as a subgroup of index $2$, via the assignment $(x,y,z) = (ab,bc,ca)$. 
Full triangle groups allow the consideration of reflections (orientation-reversing elements 
of order $2$) in many of the contexts mentioned previously. 

Next, we turn to the matter of density. 

\begin{definition}[Natural density]
Let $S$ be a subset of the positive integers $\N$. Define $d_x(S)$ by

$$d_x(S)=\frac{|S\cap\{1,\dots,x\}|}{x}$$

and, define the natural density $d(S)$ of $S$ to be $d(S)=\lim_{x\rightarrow\infty} d_x(S)$ when the limit exists.

\end{definition}

In some sense, $d(S)$ tells us the probability that a uniformly randomly chosen positive integer is an element of $S$, 
or what proportion of the natural numbers is contained in $S$. 
For example, $d(S) = 0$ when $S$ is finite, and we would expect that a random natural number 
is divisible by a given $k \in \N$ about $1/k$ of the time -- indeed
$$
d(k\mathbb{N})=\lim_{n \rightarrow\infty}\frac{|k\mathbb{N}\cap \{1,\dots,n\}|}{|\{1,\dots,n\}|} 
= \lim_{n\rightarrow\infty}\,\frac{\left\lfloor\frac{n}{k}\right\rfloor}{n} = \frac{1}{k}.
$$

In this paper, we answer the following: 

\begin{question*}
Let $\mathcal{Q}(r,s,t)$ be the set of orders of the finite quotients of the ordinary triangle group $\Delta^+(r,s,t)$. 
What is the natural density of $\mathcal{Q}(r,s,t)$? 
\end{question*}

It seems that the origin of this question was some work by Michael Larsen \cite{Larsen}, who was interested 
in the natural density of the orders of finite quotients of $\Delta^+(2,3,7)$ in order to deduce how 
often the Hurwitz bound (mentioned earlier) is attained. Larsen used the classification of finite simple groups 
and the notion of a `quotient dimension' of a linear algebraic group to prove that this density is zero, 
not only for the case $(r,s,t) = (2,3,7)$ but also for the six cases where $\frac{11}{12} < \frac{1}{r}+\frac{1}{s}+\frac{1}{t} < 1$. In fact, for these cases, Larsen proves the stronger fact that $|\mathcal{Q}(r,s,t)\cap \{1,\dots,x\}|\sim x^{1/q}$ where $q\in\{2,3\}$.

Independently (and some years later), May and Zimmerman \cite{MZ}, showed that the set of {\em odd\/}  orders of finite quotients of $\Delta^+(3,3,t)$ has density 0 when $t = 5, 7, 9$ and $11$. This was taken further in \cite{Tucker} by Tom Tucker, who recognised the importance of large cyclic normal subgroups in the quotients. By using a 1976 theorem of Bertram \cite{Bertram} on such groups (advised to him by my PhD supervisor Marston Conder), he proved that if any one of $r$, $s$ and $t$ is coprime to the other two, then $d(\mathcal{Q}(r,s,t))=0$, and asked how this might be extended.
In particular, we note that Tucker's application of Bertram's theorem showed that  $d(\mathcal{Q}(2,3,7))=0$ but did not require the classification of finite simple groups. 

By considering how and why Tucker's approach worked, it became clear that a modification of Bertram's theorem might help extend the answer to the above question to a much larger family of triples $(r,s,t)$, but this required a significant extension of the asymptotic number theory used in \cite{Bertram}. Posting a question \cite{MQ} on {\rm math}{\em overflow} (as suggested by my PhD co-supervisor Gabriel Verret) led me to a very helpful interpretation by Terrence Tao \cite{Tao} of the Tur{\' a}n-Kubilius Inequality (proved in general in the 1950s by Kubilius \cite{Kubilyus}, following a special case in the 1930s by Tur{\' a}n \cite{Turan} in the 1930s), and consequently to the main theorem of this paper: 

\begin{theorem}
\label{thm:main}
For every triple $(r,s,t)$ of positive integers, the natural density of the orders of the finite 
quotients of the ordinary triangle group $\Delta^+(r,s,t)$ is zero.
\end{theorem}

Note that this theorem covers all three kinds of ordinary triangle groups. 
It is trivial for the spherical case (in which the groups are finite), and is also quite easy for the toroidal case, 
because up to isomorphism the groups are $\Delta^+(2,3,6)$, $\Delta^+(2,4,4)$ and $\Delta^+(3,3,3)$, 
with the third of those being isomorphic to a subgroup of index $2$ in the first, 
and by observations by Coxeter and Moser in \cite{CoxeterMoser}, smooth finite quotients of the first two 
all have orders of the form $6(b^2 + bc + c^2)$ and $4(b^2+c^2)$ respectively, and hence density zero.
The hyperbolic case is much more difficult, and accordingly, for the rest of this paper, every 
ordinary triangle group can be assumed to be hyperbolic.

Our proof essentially involves using the Tur{\' a}n-Kubilius Inequality and elements of Bertram's work 
in order to construct for each triple $(r,s,t)$ a set of positive integers with natural density $1$ none of which can be the order of a quotient of the ordinary triangle group $\Delta^+(r,s,t)$.

Also Theorem \ref{thm:main} has the following immediate consequence for the full triangle groups:

\begin{corollary}
\label{cor:main2}
For every triple $(r,s,t)$ of positive integers, the natural density of the orders of the finite 
quotients of the full triangle group $\Delta(r,s,t)$ is zero.
\end{corollary}

In section \ref{sec:Background}, we give some further helpful background (involving group theory
and number theory, including some observations by Bertram in \cite{Bertram}).  
In Section \ref{sec:TK} we present the Tur{\' a}n-Kubilius inequality, and give an application 
which will be important later in the paper. Then in section \ref{sec:ProofMainThm}, we prove our main theorem.

In a sequel to this paper, (by Marston Conder and Gabriel Verret and myself) it will be shown that will prove that the natural density of the  orders of finite quotients of the free product $C_r * C_s = \langle\, x,y \ | \ x^r = y^s = 1 \,\rangle$ 
is zero whenever at least one of $r$ and $r$ is odd, but positive whenever both $r$ and $s$ are even, 
and use this to prove that the natural density of the orders of finite edge-transitive $3$-valent graphs is zero, 
together with similar consequences for arc-transitive graphs of higher valency. 

\section{Further background}
\label{sec:Background}

First, we give some elementary properties of quotients of ordinary triangle groups, 
which will provide a taste of what comes later.

\begin{proposition}
\label{prop:threeprops}
${}$ \\[-20pt] 
\begin{enumerate}[{\rm (a)}]
\item Let $G$ be a finite quotient of $\Delta^+(r,s,t)$ of order $n$ with a cyclic normal Sylow $p$-subgroup $P$, 
such that $C_G(P)$ has index $1$ or $2$ in $G$. Then $p$ divides at least one of $r,s$ and $t$.  \\[-16pt] 
\item Let $n$ be a positive integer having a prime factor $p$ such that
{\rm (i)} $p$ is coprime to $n/p$, and {\rm (ii)} the only positive integer divisor of $n$ congruent to $1$ mod $p$ is $1$. 
Then if at least two of $r$, $s$ and $t$ are coprime to $\frac{p-1}{2}$ and each of $r$, $s$ and $t$ is coprime to $p$, then $n$ cannot be the order of a finite quotient of $\Delta^+(r,s,t)$.
\\[-16pt] 
\item Let $G$ be a finite group with a normal Sylow $p$-subgroup of order $p$, 
such that $G$ is generated by two non-trivial elements $u$ and $v$ having orders $k$ and $\ell$, both coprime to $p$. 
Then {\rm (i)} $\gcd(k,p\!-\!1) \ne 1$ and  $\gcd(\ell,p\!-\!1) \ne 1$, and {\rm (ii)} if $\gcd(k, p\!-\!1) = \gcd(\ell, p\!-\!1) = 2$, 
then $p$ divides the order of $uv$. 
\end{enumerate}
\end{proposition}

\begin{proof}
For part (a), the Schur-Zassenhaus Theorem tells us that $G$ is a semi-direct product $P \rtimes H$ 
of the normal subgroup $P$ by some complementary $p'$-subgroup $H$.  
If $C_G(P)$ has index $1$ in $G$, then $P$ is central in $G$, so $G$ is a direct product of $H$ and $P$, 
which implies that $P$ is a cyclic $p$-quotient of $G$ and hence of $\Delta^+(r,s,t)$, 
and so $p$ must divide at least two of $r$, $s$ and $t$.
On the other hand, if $C_G(P)$ has index $2$ in $G$, then $C_G(P)$ is a direct product of $C_H(P)$ and $P$, 
with $C_H(P)$ being characteristic in $C_G(P)$ and hence normal in $G$, so the quotient $G/C_H(P)$ is 
isomorphic to a semi-direct product $P \rtimes C_2$, which implies that $p$ divides at least one of $r$, $s$ and $t$ 
(since $P \rtimes C_2$ cannot be generated by two elements of order~$2$ without being dihedral of order $2|P|$). 

For part (b), assume the contrary, and let $G$ be a finite quotient of $\Delta^+(r,s,t)$ of order $n$. 
Then, by Sylow theory, conditions (i) and (ii) imply that $G$ has a cyclic normal Sylow $p$-subgroup $P$ of order $p$, 
and hence that $G$ is isomorphic to a semi-direct product $P \rtimes H$ for some $p'$-subgroup $H$. 
Also $G/C_G(P)$ is isomorphic to a subgroup of $\Aut(P) \cong C_{p-1}$, and so the images in $G$ of the 
three generators $x,y$ and $z$ of $\Delta^+(r,s,t)$ either centralise $P$ or invert each element of $P$ by conjugation, 
and therefore $C_G(P)$ has index $1$ or $2$ in $G$.  
But then part (a) implies that $p$ divides at least one of $r$, $s$ and $t$, a contradiction.

Finally, part (c) can be proved using aspects of the arguments in parts (a) and (b). 
The given group $G$ is a semi-direct product $P \rtimes H$ of the cyclic normal Sylow $p$-subgroup $P$ 
by some complementary $p'$-subgroup $H$, with $N = C_H(P)$ being normal in $G$. 
Letting $\overline{J} = JN/N$ for every subgroup $J$ of $G$, 
we see that $\overline{G} \cong \overline{P} \rtimes \overline{H}$, with $\overline{H} = H/C_H(P)$ isomorphic to 
a subgroup of $\Aut(P) \cong C_{p-1}$.
Next, since $G$ is generated by a pair of elements with orders coprime to $p$, so is $\overline{G}$ 
and hence so is ${\overline H}$, and therefore ${\overline H}$ is isomorphic to a non-trivial cyclic subgroup of $\Aut(P)$, 
It follows that the images ${\overline u}$ and ${\overline v}$ of $u$ and $v$ in $\overline{G}$ are non-trivial, 
and so $\gcd(k,p\!-\!1) \ne 1$ and  $\gcd(\ell,p\!-\!1) \ne 1$, proving (i). 
Finally, if $\gcd(k,p\!-\!1) = 2$ and  $\gcd(\ell,p\!-\!1) = 2$, then ${\overline u}$ and ${\overline v}$ have order $2$, 
so $\overline{H} \cong C_2$ and therefore $\overline{G}$ is dihedral of order $2p$, 
which implies that $p$ divides the order of ${\overline u}{\overline v}$ and hence the order of $uv$. 
\end{proof}

Now we turn to some number theory that will be helpful:

\begin{definition}
Let $A$ be any set of positive integers. Then we call
$$
\sum _{n \in A}\frac{1}{n}
$$
the {\em logarithmic size\/} of $A$, and denote it by $\ell(A)$. 
If the given sum diverges, we say that $A$ has infinite logarithmic size.
\end{definition}

A well-known theorem of Dirichlet states whenever $a$ and $b$ are two coprime positive integers, 
then the arithmetic sequence $(a+bn)_{n \in \N}$ contains infinitely many primes. 
Dirichlet also proved the following stronger theorem, which can be found in \cite{Narkiewicz}: 

\begin{theorem}[Dirichlet's strong theorem]
\label{thm:StrongDirichlet}
For any given pair of coprime positive integers $a$ and $b$, the logarithmic size of the set of 
primes congruent to $a$ modulo $b$ is infinite. 
In other words, if $\mathcal{V}_x$ is the set of all primes $p\le x$ such that $p\equiv a$ mod $b$, 
then
$$
\sum_{p\in\mathcal{V}_x}\frac{1}{p}\rightarrow \infty
$$
as $x\rightarrow\infty$,  
and the logarithmic size of $\mathcal{V}_x$ is given by 
$$
\ell(\mathcal{V}_x) = \left(\frac{1}{\phi(b)}+o(1)\right)\log\log x.
$$
\end{theorem}

\begin{corollary}
\label{SDC}
Let $\mathcal{V}_x$ be as defined in Theorem {\em\ref{thm:StrongDirichlet}}, 
and let $\delta$ be any real number such that $0 < \delta < 1$.  
Then the logarithmic size of the set $\mathcal{V}'_x=\{ p \in \mathcal{V}_x \mid (\log x)^{1+\delta} < p \le x\}$
converges to infinity.  
Equivalently, 
$$
\sum_{p\in\mathcal{V}_x'}\frac{1}{p}\rightarrow \infty
    \quad \hbox{ as } \ x\rightarrow\infty.
$$
\end{corollary}

\begin{proof}
${}$\\[-33pt] 
\begin{align*}
   \ell(\mathcal{V}'_x) 
   &= \ell(\mathcal{V}_x\cap ((\log x)^{1+\delta},x]) \\[+3pt]
   &= \ell(\mathcal{V}_x) - \ell(\mathcal{V}_{(\log x)^{1+\delta}}) \\
   &= \left(\frac{1}{\phi(b)}+o(1)\right)\log(\log x) - \left(\frac{1}{\phi(b)}+o(1)\right)\log(\log((\log x)^{1+\delta})) \\
   &=  \left(\frac{1}{\phi(b)}+o(1)\right)\log\left(\frac{\log x}{\log(\log x)^{1+\delta}}\right), 
       \,\hbox{ which converges to infinity.} 
\end{align*}
${}$ \\[-32pt] 
\end{proof}

\medskip
Next, we outline some key constituents of Bertram's approach, 
namely three observations made as Lemmas \ref{lem:B1} to \ref{lem:B3} in \cite{Bertram}, 
with each of $f$, $g$ and $h$ being a function from $\R^+$ to $\R^{+}\!.$ 
One of these needs the following:

\begin{definition}
Let $x\in\N$ and let ${\cal P}(x)$ be some number-theoretic property possibly dependent on $x$. Then let $S_x$ be the set of all $n \le x$ in $\N$ such that $n$ satisfies ${\cal P}(x)$. Then we say that {\em almost all positive integers $\le x$ satisfy ${\cal P}(x)$\/} if $\,\lim_{x\rightarrow \infty}  d_x(S_x) = 1$ or equivalently if $\lim_{x\rightarrow\infty}\frac{|S_x|}{x}=1$.
\end{definition}

For example, if $S_x$ is the set of positive integers less than $x$ but greater than $\log x$, we would say that almost all positive integers $\le x$ are greater than $\log x$ because $\lim_{x\rightarrow\infty}\frac{|S_x|}{x}=1$.
\begin{observation*}
    Note that if almost all positive integers $\le x$ satisfy ${\cal P}(x)$, then the set of positive integers satisfying ${\cal P}(x)$ for at least one $x$ has natural density $1$.

    $$
    d_x\left(\bigcup_{y\in\N}S_y\right) = \frac{|\bigcup_{y\in\N}S_y \cap \{1,\dots,x\}|}{x} \ge \frac{|S_x\cap\{1,\dots,x\}|}{x} = \frac{|S_x|}{x}\rightarrow 1.
    $$

\end{observation*}

\begin{lemma}
\label{lem:B1}
For all $x \in \R^+,$ the number of positive integers $n\le x$ such that $p^2$ divides $n$ for some prime $p>f(x)$ 
is less than $\frac{x}{f(x)}$.
\end{lemma}

\begin{lemma}
\label{lem:B2}
The number of positive integers $n\le x$ having a prime factor $p>f(x)$, and simultaneously having a divisor $d>1$ satisfying $d\equiv 1$ mod $p$, is less than $\frac{x((\log x) +1)}{f(x)}$.
\end{lemma}

\begin{lemma}
\label{lem:B3}
The number of positive integers $n\le x$ which have a divisor $d\ge h(x)$, such that each prime factor of $d$ is less than $g(x)$, is less than $\frac{x(\log(g(x))+c)}{\log(h(x))}$ for some constant $c$.
\end{lemma}

Proofs of the three lemmas above can be found in \cite[pp.\,64--65]{Bertram}.

\section{The Tur{\' a}n-Kubilius Inequality}
\label{sec:TK}

Now we describe and apply the Tur{\' a}n-Kubilius inequality. 
Tur{\' a}n first proved a special case of this inequality  in order to simplify the proof of the famous Hardy-Ramanujan theorem 
(see \cite{Turan}), and Tur{\' a}n's work was later reformulated and generalised by Kubilius (see \cite{Kubilyus}, 
or \cite{Kubilius} for a version in English).  

As is customary in number theory, here a complex-valued  arithmetic function $f\!: \Z \to \C$ is said to be {\em additive\/}  
if $f(mn)=f(m)+f(n)$ whenever $m$ and $n$ are relatively prime.

\begin{theorem}[Turan-Kubilius]\
\label{thm:TK} 
Let $f\!: \Z \to \C$ be an additive complex-valued arithmetic function, and write $p$ for an arbitrary prime,
and $\nu$ for an arbitrary positive integer. 
Also, for given $x \in \N$, let $\mathcal{Z}_x$ be the set of prime powers less than or equal to $x$, and write
$$
A(x)=\sum_{p^\nu \in \mathcal{Z}_x}\frac{f(p^\nu)}{p^\nu}\,(1-\frac{1}{p})
 \quad \hbox{ and } \quad
B(x)^2=\sum_{p^\nu\in\mathcal{Z}_x}\frac{|f(p^\nu)|^2}{p^\nu}.
$$
Then there exists a function $g\!: \Z \to \R$ which converges to zero, such that 
$$
\frac{1}{x}\sum_{n\le x}|f(n)-A(x)|^2 \le (2+g(x))B(x)^2 \ \ \hbox{ for all } x\ge 2.
$$
\end{theorem}

\begin{proof}
This can be found in a book by Tenenbaum \cite{Tenenbaum} and/or one by Narkiewicz \cite{Narkiewicz}.
\end{proof}

Next, we prove two things that lead to an important application. 

\begin{lemma}
\label{lem:Infntsize}
Let $m$ be an integer greater than $1$, and let $\mathcal{P}$ be the set of all primes $p$ 
such that $\frac{p-1}{2}$ is coprime to $m$. 
Then $\mathcal{P}$ contains all of the prime terms (perhaps excluding $2$) of some arithmetic sequence $(a+bn)_{n\ge 1}$ 
where $\gcd(a,b) = 1$, and in particular, contains infinitely many such primes.
\end{lemma}

\begin{proof}
Let $\mathcal{Q}$ be the arithmetic sequence $\{2m-1+2mn\}_{n\ge 2}$. Clearly, $2m-1$ and $2m$ are coprime, so $\mathcal{Q}$ contains infinitely many primes. If a prime $p>2$ is in $\mathcal{Q}$ then $p=2m-1+2mn$ for some $n$, hence $\frac{p-1}{2}=(n+1)m-1$ and so $\frac{p-1}{2}$ is coprime to $m$, with that we are done.
\end{proof}

\begin{corollary}
\label{cor:logsizePx}
Let $m$ and $\mathcal{P}$ be as defined in Lemma {\em \ref{lem:Infntsize}}. 
Then for any $\delta \in \R$ such that $0 < \delta < 1$, 
the logarithmic size of $\Px=\mathcal{P}\cap ((\log x)^{1+\delta},\dots,x]$ converges to infinity. 
\end{corollary}

\begin{proof}
By Lemma \ref{lem:Infntsize}, there exists an arithmetic sequence $(a+bn)_{n\ge 1}$ with $\gcd(a,b) = 1$ 
that contains an infinite set $P(a,b)$ of prime terms.
Then $P(a,b)_x = P(a,b)\cap ((\log x)^{1+\delta},\dots,x]\subset \Px$ and so $\ell(P(a,b)_x)\le \ell(\Px)$ for all $x$, 
and then since $\ell(P(a,b)_x)$ converges to infinity (by Corollary \ref{SDC}), also $\ell(\Px)$ converges to infinity.
\end{proof}

\begin{proposition}
\label{prop:ATK}
Let $m$ be an integer greater than $1$, let $\delta \in \R$ with $0 < \delta < 1$, 
and for $x \in \N$, let $S_x$ be the set of positive integers $n\le x$ such that there exists a prime factor $p$ of $n$ 
with $\frac{p-1}{2}$ coprime to $m$ and $p > (\log x)^{1+\delta}$.
Then $d_x(S_x)$ converges to $1$ as $x$ tends to infinity.
\end{proposition}

\begin{proof}
Let $\Px$ denote the set of primes $p$ such that $\gcd(p,m) = 1$ and $(\log x)^{1+\delta}< p\le x$.
We wish to apply the Turan-Kubilius Inequality using an additive function $f$ with the property 
that if $f(n)>1$, then $n\in S_x$, so we take $f$ to be the additive function for which $f(n)$ is the number 
of prime factors of $n$ that lie in $\Px$.
Note that for $\nu \in  \N$ we have $f(p^\nu)=1$ when $p\in\mathcal{P}_x$, and $f(p^\nu)=0$ otherwise. 

Next, let $\mathcal{Z}^\Px$ be the set of positive integers less than or equal to $x$ which are powers (including first powers) of primes in $P_x$. Then, we may write
$$
A(x) = \sum_{p^\nu \in \mathcal{Z}_x}\frac{f(p^\nu)}{p^\nu}\,(1-\frac{1}{p}) 
    = \sum_{p^\nu \in\mathcal{Z}^\mathcal{P}_x}\,\frac{1}{p^\nu}\,(1-\frac{1}{p})
\quad \hbox{ and } \quad  
B(x)^2  =  \sum_{p^\nu\in\mathcal{Z}_x}\frac{|f(p^\nu)|^2}{p^\nu} \\
     =  \sum_{p^\nu\in\mathcal{Z}^\mathcal{P}_x}\,\frac{1}{p^\nu},    
$$
and so by Theorem \ref{thm:TK}, 
$$
\frac{1}{x}\,\sum_{n\le x}\,|f(n)-A(x)|^2 \,\le\, (2+g(x))B(x)^2 \quad \hbox{for all } x \ge 2.
$$

Next, let $\epsilon$ be any real constant $> 0$, and say that $n\in\mathbb{N}$ is `good' if $n\le x$ and $|f(n)-A(x)|< B(x)^{1+\epsilon}$, or `bad' if $n\le x$ and $|f(n)-A(x)|\ge B(x)^{1+\epsilon}$. 
Letting $n_b$ be the number of bad $n\le x$, we see that  

\begin{equation}
 \begin{split}
            \frac{1}{x}\sum_{n\le x}|f(n)-A(x)|^2 &= \frac{1}{x}\sum_{\text{good}\ n}|f(n)-A(x)|^2+\frac{1}{x}\sum_{\text{bad}\ n}|f(n)-A(x)|^2\\
                 &\ge \frac{1}{x}\sum_{\text{good}\ n}|f(n)-A(x)|^2 + \frac{1}{x}\sum_{\text{bad}\ n}(B(x)^{1+\epsilon})^2\\
                 &= \frac{1}{x}\sum_{\text{good}\ n}|f(n)-A(x)|^2 + \frac{n_bB(x)^{2+2\epsilon}}{x}\\
                 &\ge \frac{n_bB(x)^{2+2\epsilon}}{x}, 
 \end{split}
\end{equation}
and hence that 
$$
\frac{n_bB(x)^{2+2\epsilon}}{x}\le\frac{1}{x}\sum_{n\le x}|f(n)-A(x)|^2 \le (2+g(x))B(x)^2,
$$
and therefore 
\begin{equation}\label{nb}
\frac{n_b}{x} \le \frac{2+g(x)}{B(x)^{2\epsilon}}.
\end{equation}

Note that $B(x)^2$ is strictly greater than $\ell(\Px)=\sum_{p\in\Px}\frac{1}{p}$, since $B(x)^2$ includes not only reciprocals of
 primes but also reciprocals of prime-powers. 
Also, Corollary \ref{cor:logsizePx} shows that $\ell(\Px)$ converges to infinity, and therefore so does $B(x)^2$. 
By inequality (\ref{nb}), this implies that the density of the `bad' positive integers is zero, 
so almost all positive integers are `good', and hence that  $|f(n)-A(x)|< 
B(x)^{1+\epsilon}$ almost always. Intuitively this says that if $n$ is good then 
$f(n)$ is at most $B(x)^{1+\epsilon}$ steps away from $A(x)$ when $n\le x$. But as 
$A(x)$ should converge to infinity, and should be large relative to 
$B(x)^{1+\epsilon}$, this also requires $f(n)$ to be large when $n$ is good, 
which, as we have shown, is almost always the case.

It remains to prove that $A(x)$ converges to infinity 
and is large relative to $B(x)^{1+\epsilon}$. 
In fact, all we need to do is show that $A(x)-B(x)^{1+\epsilon}\ge 1$ for sufficiently large $x$. 
We may start by noting that
$$ A(x) 
\ = \ \sum_{p^\nu \in\mathcal{Z}^\mathcal{P}_x}\,\frac{1}{p^\nu}\,(1-\frac{1}{p})
\ = \ \sum_{p^\nu\in \mathcal{Z}^\mathcal{P}_x}\frac{1}{p^\nu}-\sum_{p^\nu\in \mathcal{Z}^\mathcal{P}_x}\frac{1}{p^{v+1}} 
\ = \ B(x)^2-\sum_{p^\nu\in \mathcal{Z}^\mathcal{P}_x}\frac{1}{p^{v+1}}.    
$$

We now seek to bound the sum $G(x)=\sum_{p^\nu\in \mathcal{Z}^\mathcal{P}_x}\frac{1}{p^{v+1}}$. By simple algebraic manipulation (involving the geometric series $\sum_{\nu = 0}^{\infty}(\frac{1}{p})^\nu = \frac{1}{1-\frac{1}{p}}$ 
and the telescoping series $\sum_{n=2}^{\infty} \frac{1}{n^2-n} \ = \sum_{n=2}^{\infty} (\frac{1}{n-1}-\frac{1}{n}) = 1$), 
we see that
\begin{align*}
    G(x)& \ = \ \sum_{p^\nu\in \mathcal{Z}^\mathcal{P}_x}\frac{1}{p^{v+1}}= \sum_{p\in \Px}\sum_{\nu = 2}^{\lfloor\log_p(x)\rfloor}\frac{1}{p^\nu}\\
        &\ < \ \sum_{p\in \Px}\, \sum_{\nu = 2}^{\infty}\,\frac{1}{p^\nu} \ = \ \sum_{p\in \Px}\left(\left(\sum_{\nu = 0}^{\infty}\frac{1}{p^\nu}\right)-1-\frac{1}{p}\right) \ = \ \sum_{p\in \Px}\left( \frac{1}{1-\frac{1}{p}}-1 -\frac{1}{p} \right)\\
        &\ = \ \sum_{p\in \Px}\frac{p^2 -p(p-1)-(p-1)}{p^2-p}\ = \ \sum_{p\in \Px}\frac{1}{p^2-p} \ < \ \sum_{n=2}^{\infty} \frac{1}{n^2-n} \ = \ 1.
\end{align*}

Thus $0< G(x)< 1$, and it follows that $B(x)^2-1 < B(x)^2 - G(x) = A(x) < B(x)^2$. 

Then since $B(x)^2$ converges to infinity, so does $A(x)$, 
and moreover, there exists large enough $x_0 \in \mathbb{N}$ and small enough $\epsilon>0$ to ensure that $B(x)^2-1>B(x)^{1+\epsilon}+1$ for all $x\ge x_0$. 
Hence, for large enough $x$, we have 
\begin{equation}\label{ineq}
    A(x)-B(x)^{1+\epsilon}>1.
\end{equation}

Next, recall the inequality $|f(n)-A(x)|<B(x)^{1+\epsilon}$ derived in the paragraph following (\ref{nb}). 
This inequality holds for almost all $n$, and by expanding it and combining this with (\ref{ineq}), we obtain
$$
1 < A(x)-B(x)^{1+\epsilon}<f(n)<A(x)+B(x)^{1+\epsilon},
$$
which again holds for almost all positive integers $n \le x$. 
Hence in particular, $f(n)>1$, and so $n\in S_x$. 

It follows that the natural density of $S_x$ approaches $1$ as $x\rightarrow\infty$. 

Furthermore, if $n\not\in S_x$, then $f(n)=0$, and so $|f(n)-A(x)|=|A(x)|\ge B(x)^{1+\epsilon}$ for large enough $x$, 
and hence $n$ is bad, which by (\ref{nb})) immediately implies that $$|\{1,\dots,x\}\setminus S_x|\le n_b \le x\left( \frac{2+g(x)}{B(x)^{2\epsilon}}\right).$$
${}$\\[-33pt] 
\end{proof}

\section{Proof of our main theorem}
\label{sec:ProofMainThm}

We now have enough to prove Theorem \ref{thm:main}, namely that for every triple $(r,s,t)$ of positive integers, the natural density of the orders of the finite quotients of the ordinary triangle group $\Delta^+(r,s,t)$ is zero.

\begin{proof}
Let $S_x$ be the same set as in proposition \ref{prop:ATK}. Namely, fixing some real number $0<\delta<1$ and natural number $m>1$ then $S_x$ is the set of natural numbers $n<x$ with a prime factor $p>(\log x)^{1+\delta}$ such that $\gcd(\frac{p-1}{2},m)=1$.

Let $H_x$ denote the set of natural numbers $n$ less than or equal to $x$ with the property that for all prime factors $p>(\log x)^{1+\delta}$ of $n$ then $p^2$ is not a factor of $n$, and let $G_x$ denote the set of natural numbers less than or equal to $x$ with the property that for any $n\in G_x$ if $p>(\log x)^{1+\delta}$ is a prime factor of $n$ then there is no factor $d>1$ of $n$ such that $d\equiv 1 \hbox{ mod }  p$.

Next, let $K_x=S_x\cap G_x\cap H_x$. So any natural number $n\in K_x$ has a prime factor $p$ such that 

\begin{itemize}
    \item $p>(\log x)^{1+\delta}$,
    \item $p^2$ does not divide $n$,
    \item $\gcd(\frac{p-1}{2},m)$=1,
    \item for all factors $d\ne 1$ of $n$, then $d\not\equiv 1 \hbox{ mod }p$.
\end{itemize}
Fix three natural numbers $r,s,t>1$ and set $m=rst$ and choose $x$ large enough to ensure that $(\log x)^{1+\delta}>rst$, then by proposition \ref{prop:threeprops} (b), no element $n$ of $K_x$ can be the order of a finite quotient of $\Delta^+(r,s,t)$. Hence, $\mathcal{Q}(r,s,t)\cap K_x=\emptyset$ meaning $\mathcal{Q}(r,s,t)\cap\{1,\dots,x\}\subset \{1,\dots,x\}\setminus K_x$. However by proposition \ref{prop:ATK}, lemma \ref{lem:B1}, and lemma \ref{lem:B2} it must be that if $\epsilon>0$ is sufficiently small, then $|\{1,\dots,x\}\setminus K_x|$ can be no larger than

$$
N(x) = \frac{x}{(\log x)^{1+\delta}} + \frac{x((\log x) +1)}{(\log x)^{1+\delta}} + \frac{x(2+g(x))}{B(x)^{2\epsilon}}.
$$

But, clearly $N(x)/x \rightarrow 0$ as $x\rightarrow\infty$, and hence the natural density of $\mathcal{Q}(r,s,t)$ must be zero, this completes the proof of our main theorem.

\end{proof}

\bigskip

{\Large\bf Acknowledgements}
\medskip

The author is grateful to his PhD supervisor, Marston Conder and co-supervisors Jeroen Schillewaert and Gabriel Verret,  
and to New Zealand's Marsden Fund for its financial support for his PhD studies via Marston Conder's grant UOA2030. He also acknowledges the helpful use of the {\sc Magma} system \cite{Magma} in investigating quotients of triangle groups in some particular cases. Finally, he is grateful to Ben McReynolds for informing him about the paper by Larsen \cite{Larsen} after the first version of the current paper was completed.


\bigskip

\end{document}